a father age is added to his son age to be 69 years. in three years time the father age will be 4times old than his son. find their age?

Answer by Boreal(15235)

father = x
son=y
x+y=69
(x+3)=4(y+3); in 3 years
x+3=4y+12
x=4y+9
substitute into the first equation
4y+9+y=69
5y+9=69
5y=60
y=12 (son)
x=57 (father)
in 3 years, 60 and 15, which is 4 times.